# Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models

### v1

@@ 3.3.3 Key Takeaways on the Jailbreak Process
Based on the above analysis, we summarize the jailbreak process as follows: (1) Jailbreak takes effect in the low layers, contrary to prior work Zhou et al. (2024), which suggested it begins in the middle layers. (2) As the attack progresses, jailbreak continues to operate, shifting the activation outside the safety boundary and causing the model to be deceived. (3) When comparing all layers, the low and the middle layers exhibit the strongest shift and smallest safety boundaries, demonstrating their importance in the jailbreak process.
@@ 4 Activation Boundary Defense
Based on the above findings, we propose a defense called Activation Boundary Defense (ABD), whose core idea is to confine the activations within the safety boundaries. The workflow of ABD is shown in Figure 4. ABD includes a penalty function and a Bayesian optimization process. The penalty constrains activations within safety boundaries, while Bayesian optimization iteratively adjusts the affected layers and the penalty parameters.
@@ 4.1 Penalty Function
We empirically find that the activation distributions in each layer can be approximated by a normal distribution. The proof is as follows. For each layer $l$, we examine two distributions: the activation coordinate distribution $\mathcal{D}^{l}(x)$ and a normal distribution $\mathcal{D_{N}}^{l}(x)$ with the same mean $\mu^{l}_{\mathcal{D}}$ and standard deviation $\sigma^{l}_{\mathcal{D}}$. To measure their similarity, we compute the Jensen-Shannon (JS) divergence Lin (1991) between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$. Across all layers, the maximum JS divergence is 0.0839, and the mean is 0.0575, and both are well below 0.1. These low JS divergence values indicate strong similarity between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$, which supports the validity of approximating $\mathcal{D}^{l}(x)$ with $\mathcal{D_{N}}^{l}(x)$.
@@
where $\alpha^{l}\geq 0$ and $\beta^{l}\geq 0$ are hyperparameters. Visualization is given in Appendix D.1.
This function meets the stated expectations. It is symmetric about the mean $\mu^{l}_{\mathcal{D}}$, ensuring fair penalization for deviations above and below this central value. The penalty strength $|x-x^{\prime}|$ increases nonlinearly with the distance from $\mu^{l}_{\mathcal{D}}$ because $\mathrm{tanh(\cdot)}$ amplifies more significant deviations with its steep slope, reflecting stronger penalization for outliers. Moreover, the function mostly penalizes outliers. Within the range $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}],x^{\prime}\approx x$, resulting in a negligible penalty for values close to the mean. The hyperparameters $\alpha^{l}$ and $\beta^{l}$ govern the behavior of the penalty function. The parameter $\alpha^{l}$ determines the maximum range of $x^{\prime}$, ensuring all coordinates are constrained within $(-\alpha^{l}+\mu^{l}_{\mathcal{D}},\alpha^{l}+\mu^{l}_{\mathcal{D}})$ after penalty application. Meanwhile, $\beta^{l}$ controls the size of the unmodified region $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}]$; increasing $\beta^{l}$ expands this region, while decreasing $\beta^{l}$ narrows it.
@@ 4.2 Bayesian Optimization-Based Tuning
Model Jailbreak No Defense Paraphrase PPL Retokenization SafeDecoding Self-Exam Self-Reminder ABD (Ours)
Vicuna No attack 88% 82% 90% 66% 88% 100% 100% 100%
GCG-Individual 4% 86% 78% 62% 100% 86% 100% 100%
AutoDAN 6% 26% 12% 36% 78% 30% 30% 44%
PAIR 30% 62% 40% 30% 94% 86% 100% 76%
DeepInception 10% 2% 2% 0% 98% 18% 40% 58%
Llama-2 No attack 100% 100% 100% 94% 100% 100% 100% 100%
GCG-Individual 50% 98% 100% 98% 100% 64% 100% 100%
AutoDAN 98% 94% 98% 94% 100% 100% 98% 100%
PAIR 72% 90% 82% 78% 92% 100% 86% 92%
DeepInception 74% 82% 88% 60% 100% 94% 96% 100%
Table 1: The Defense Success Rate (DSR) of different defense methods across Vicuna-7B-v1.3 (simplified as Vicuna) and Llama-2-7B-chat (simplified as Llama-2) along with different attack types, where a higher DSR indicates better defense effectiveness. For ABD, the best and second performance across all defenses are highlighted.
Another crucial aspect of our method is the configuration of the hyperparameters and the application scope of the penalty function. Specifically, we must (1) minimize the number of affected activation coordinates to limit unnecessary perturbations, and 2) for each affected layer $l$, find the values of $\alpha^{l}$ and $\beta^{l}$ that best improve the model’s resilience against adversarial attacks. To accomplish these goals, we use a Bayesian Optimization (BO)-based Jones et al. (1998); Mockus (1974) tuning method. Concretely, we define two core objectives.
@@ 5 Experiments
Table 2: Comparison of defenses on Runtime per Query and Just-Eval metrics in Vicuna-7B-v1.3. $\downarrow$: smaller is better; $\uparrow$: larger is better. For ABD, the best and second performances are highlighted. ABD preserves the model’s general performance, adding less than 0.1 seconds to runtime while producing high-quality outputs with leading evaluation scores. It has the best overall performance across all defenses.
@@ 5.1 Settings
We consider two widely used open-source LLMs as the target models in our experiments: Llama-2-7B-Chat Touvron et al. (2023) and Vicuna-7B-v1.3. Llama-2-7B-Chat is specially trained for safety alignment, while Vicuna-7B-v1.3 is tuned based on Llama without extra alignment for safety. The configurations of the LLMs are shown in Appendix E.1.
@@
For jailbreak methods, we considered four widely applied ones: optimization-based methods include GCG-Individual and AutoDAN, which iteratively optimize a jailbreak prompt towards the goal of generating affirmative responses. Model-based jailbreak, i.e., PAIR, uses an attacker LLM to refine the jailbreak prompts. The rule-based method includes DeepInception Li et al. (2023), which crafts jailbreak prompts based on a stealthy template. Correspondingly, we applied different defense methods to the models to compare them to our model for combating jailbreak methods: Paraphrase Jain et al. (2023), Retokenization Jain et al. (2023) and Self-Reminder Xie et al. (2023) reformulate the input to avoid attack; PPL Alon and Kamfonas (2023) and Self-ExamPhute et al. (2023) defense LLMs by double-checking their outputs; SafeDecoding Xu et al. (2024) uses a tuned model to modify the output probability distribution.
@@
Following Xu et al. (2024) and Chao et al. (2023), we utilized 50 samples from AdvBench Zou et al. (2023) as the test set. We used DSR when evaluating defense effectiveness. Furthermore, following Xu et al. (2024), we adopted Just-Eval Lin et al. (2023) to measure the general abilities of LLMs before and after applying defense. Just-Eval is a comprehensive benchmark containing 1,000 diverse instructions, covering seven task types (e.g., reasoning, math, coding, etc.) and seven topics (e.g., ethics, nature, STEM, etc.). Following Lin et al. (2023), we leveraged GPT-4o-mini OpenAI (2024) to score the quality of the outputs, ranging from 1 to 5, across five aspects: helpfulness, clarity, factuality, depth, and engagement. We reported the average score for each aspect and the overall average score for different aspects, denoted as Avg. To evaluate the efficiency, we calculated the Runtime per Query for each defense, which represents the average time taken to process a single query during inference. This metric provides a direct measurement of the efficiency of defenses in real-world applications. We also utilized the above two metrics to derive an overall score that simultaneously reflects LLMs’ responding speed and quality. The calculation is shown in Appendix E.4.1.
@@
We randomly filtered 400 non-overlapping AdvBench samples to compute $\mu_{\mathcal{D}}^{l}$ and used them as a validation set. We used GCG-Universal Zou et al. (2023) to attack the validation set. GCG-Universal finds a shared jailbreak suffix for all harmful prompts that can deceive LLMs. By default, we set $w=0.7$ when calculating $\mathcal{L}_{\mathrm{Robust}}(\Theta,M)$ to slightly prioritize improving DSR while balancing defensive capability and perturbation to LLMs. Appendix D.2 presents more details about ABD optimization.
@@ 5.2 Experimental Results
In both models, low and middle layers, such as layers 5 and 12 in Vicuna-7B-v1.3 and layers 2 and 12 in Llama-2-7B-chat, are penalized, which aligns with our observed findings and proposed explanation that low and middle layers are more vulnerable to jailbreak attacks due to noticeable activation shifts.
@@
Figure 3(c) shows the percentage of jailbreak activations within the harmful activation space, namely the inclusion ratio. Before applying ABD, the inclusion ratio remains below 0.4 for various jailbreak activations, indicating that jailbreak activations lie outside of the safety boundary. After applying ABD, the ratio increases to 1, demonstrating that our method effectively constrains the jailbreak activations within the safety boundary. Additional details are provided in Appendix C.2. From a visualization perspective, Figure 1 projects activations for different prompts. Under our defense, jailbreak activations are progressively constrained within the harmful activation space under ABD, as shown in Figure 1. further confirming the effectiveness of ABD for mitigating jailbreak effects.
Statistic effectiveness of ABD.
We report the statistical defense results of ABD on the test set in Table 1. For Vicuna-7B-v1.3, which lacks specific safety alignment, the defense poses a more challenging task. Nevertheless, ABD demonstrates competitive performance. For jailbreak methods such as PAIR and DeepInception, ABD achieves a higher DSR (58%) compared to Paraphrase (2%) and Retokenization (0%), both of which require costly prompt reformulation. Against the GCG-Individual attack, ABD successfully defends against all jailbreak samples. For the well-aligned model Llama-2-7B-chat, ABD achieves 100% DSR under most jailbreak methods.
ABD is efficient and reliable.
Table 2 shows Runtime per Query, Just-Eval scores, and overall scores of Vicuna-7B-v1.3 with different defenses applied. We find that ABD has the greatest overall score across all defenses. Moreover, ABD only adds marginal extra time cost and general ability affection. Specifically, it only causes less than $1\%$ delay in each sample and less than $2\%$ perturbation in overall ability, compared to costly methods such as Paraphrase and Self-Exam. Furthermore, with ABD applied, the helpfulness, actuality, depth, and engagement also show a slight increase. We further discover that for baseline defenses such as Retokenization and Self-Reminder, LLM would generate overly simplistic outputs, which leads to smaller Runtime per Query, but they have significantly smaller Just-Eval scores.
@@ 6 Conclusion and Future Work
We conducted a comprehensive study of jailbreak mechanisms, analyzing over 30,000 samples. Our findings reveal that jailbreak shifts harmful activations outside the safety boundary in each layer, with the most severe shifts occurring in the low and middle layers. Motivated by this finding, we proposed ABD, which drives jailbreak activations back within the safety boundary, utilizing LLMs’ intrinsic sensitivity to harmful information. Our experiments suggest that ABD is both practical and efficient. In the near future, we aim to investigate the safety challenges of jailbreak attacks in multi-turn dialogue systems.
@@
For certain attack methods, under-aligned models (Vicuna-7B-v1.3) may not perform significantly as well-aligned models (Llama2-7B-chat). We believe this is because under-aligned models have unclear safety boundaries, which complicate the search for penalty functions that balance general ability and DSR. Future work could refine ABD by specifically searching for activation spaces that preserve the concept of "safety", therefore enhancing its generalizability on uncensored models.
@@
In this study, we primarily focus on single-round jailbreak scenarios. We do not extend our analysis to more complex jailbreaks that involve long contexts or multi-round dialogues, such as CFA Sun et al. (2024). As a result, the relationship between jailbreak dialogues and safety boundaries remains largely unexplored. Addressing this limitation, we plan to investigate and incorporate methods for detecting safety boundaries in multi-round dialogue scenarios as part of our future work.
@@
The aim of this research is to enhance the explainability and safety of LLMs. Our proposed jailbreak mechanism, that jailbreak shift activations out of the safety boundary, can mitigate disputes on how jailbreak happens and promote the development of both LLM explainability and safety. We highlight that the development of ABD only needs publicly available datasets and jailbreak methods and does not require designing new jailbreak methods. We demonstrate some harmful responses from LLMs only for illustration purposes. We acknowledge that ABD would cause the development of new attacks. Therefore, we will explore using random perturbation in the input sequence rather than a particular jailbreak method when optimizing to mitigate such attacks.

### v2

@@ 3.3.3 Key Takeaways on the Jailbreak Process
Based on the above analysis, we summarize the jailbreak process as follows: (1) Jailbreak takes effect in the low layers, contrary to prior work Zhou et al. (2024b), which suggested it begins in the middle layers. (2) As the attack progresses, jailbreak continues to shift the activation outside the safety boundary, causing the model to be deceived. (3) When comparing all layers, the low and the middle layers exhibit the strongest shift and smallest safety boundaries, demonstrating their importance in the jailbreak process.
@@ 4 Activation Boundary Defense
Although the safety boundary is not directly measurable, we can still use this concept to prevent jailbreaks. Here we propose a lightweight defense called Activation Boundary Defense (ABD), whose core idea is to confine the activations within the safety boundaries. The workflow of ABD is shown in Figure 4. ABD includes a penalty function and a Bayesian optimization process. The penalty constrains activations within safety boundaries, while Bayesian optimization iteratively adjusts the affected layers and the penalty parameters.
@@ 4.1 Penalty Function
We find that the activation distributions in each layer can be well approximated by a normal distribution. The proof is as follows. For each layer $l$, we examine two distributions: the activation coordinate distribution $\mathcal{D}^{l}(x)$ and a normal distribution $\mathcal{D_{N}}^{l}(x)$ with the same mean $\mu^{l}_{\mathcal{D}}$ and standard deviation $\sigma^{l}_{\mathcal{D}}$. To measure their similarity, we compute the Jensen-Shannon divergence (JS divergence) Lin (1991) between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$. Across all layers, the maximum JS divergence is 0.0839, and the mean is 0.0575, and both are well below 0.1. We visualize all JS divergences in Table 7 in Appendix E.1. These low JS divergence values indicate a strong similarity between $\mathcal{D}^{l}(x)$ and $\mathcal{D_{N}}^{l}(x)$, which supports the validity of approximating $\mathcal{D}^{l}(x)$ with $\mathcal{D_{N}}^{l}(x)$.
@@
where $\alpha^{l}\geq 0$ and $\beta^{l}\geq 0$ are hyperparameters. The penalty functions under different $\alpha^{l}$ and $\beta^{l}$ are visualized in Figure 6 in Appendix E.2.
This function meets the stated expectations. It is symmetric about the mean $\mu^{l}_{\mathcal{D}}$, ensuring fair penalization for deviations above and below this central value. The penalty strength $|x-x^{\prime}|$ increases nonlinearly with the distance from $\mu^{l}_{\mathcal{D}}$ because $\mathrm{tanh(\cdot)}$ amplifies more significant deviations with its steep slope, reflecting stronger penalization for outliers. Moreover, the function mostly penalizes outliers. Within the range $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}],x^{\prime}\approx x$, where $b^{l}$ is a relatively small number, resulting in a negligible penalty for values close to the mean. The hyperparameters $\alpha^{l}$ and $\beta^{l}$ govern the behavior of the penalty function. The parameter $\alpha^{l}$ determines the maximum range of $x^{\prime}$, ensuring all coordinates are constrained within $(-\alpha^{l}+\mu^{l}_{\mathcal{D}},\alpha^{l}+\mu^{l}_{\mathcal{D}})$ after penalty application. Meanwhile, $\beta^{l}$ controls the size of the unmodified region $[\mu^{l}_{\mathcal{D}}-b^{l},\mu^{l}_{\mathcal{D}}+b^{l}]$; increasing $\beta^{l}$ expands this region, while decreasing $\beta^{l}$ narrows it.
@@ 4.2 Bayesian Optimization-Based Tuning
Model Jailbreak No Defense Paraphrase PPL Retokenization SafeDecoding Self-Exam Self-Reminder IA ABD (Ours)
Vicuna-7B No attack 88% 82% 90% 66% 88% 100% 100% 100% 100%
GCG-Individual 4% 86% 78% 62% 100% 86% 100% 100% 100%
AutoDAN 6% 26% 12% 36% 78% 30% 30% 100% 44%
PAIR 30% 62% 40% 30% 94% 86% 100% 100% 76%
DeepInception 10% 2% 2% 0% 98% 18% 40% 100% 58%
Llama-2 No attack 100% 100% 100% 94% 100% 100% 100% 100% 100%
GCG-Individual 50% 98% 100% 98% 100% 64% 100% 100% 100%
AutoDAN 98% 94% 98% 94% 100% 100% 98% 100% 100%
PAIR 72% 90% 82% 78% 92% 100% 86% 100% 92%
DeepInception 74% 82% 88% 60% 100% 94% 96% 100% 100%
Table 1: DSR of different defense methods across Vicuna-7B and Llama-2 along with different attack types . For ABD, the best, second and third performance across all defenses are highlighted.
Model Jailbreak No Defense IA ABD (Ours)
Qwen No attack 96.00% 84.00% 98.00%
GCG-Individual 6.84% 44.64% 98.74%
Vicuna-13B No attack 98.00% 100.00% 100.00%
GCG-Individual 38.46% 100.00% 100.00%
Table 2: Representative results comparing DSR of Qwen and Vicuna-13B. The best performance is highlighted.
Another crucial aspect of ABD is determining which layers to penalize and what hyperparameters to choose. Specifically, we aim to minimize the affected layers to reduce unnecessary perturbations and optimize $\alpha^{l}$ and $\beta^{l}$ for each layer $l$ to enhance the model’s resilience against jailbreaks. To accomplish these goals, we use a Bayesian Optimization (BO)-based Jones et al. (1998) tuning method to find suitable configurations. The BO process used for tuning is detailed in Appendix E.4. Concretely, we define two core objectives.
@@ 5 Experiments
Table 3: Comparison of defenses on Runtime per Query and Just-Eval metrics in Vicuna-7B. $\downarrow$: smaller is better; $\uparrow$: larger is better. For ABD, the best and second performances among all defenses are highlighted. ABD preserves the model’s general performance, adding less than 0.1 seconds to runtime while producing high-quality outputs with leading evaluation scores. It has the best overall performance across all defenses.
@@ 5.1 Settings
We evaluated four widely used open-source LLMs of varying sizes and architectures: Llama-2-7B-Chat Touvron et al. (2023), Vicuna-7B-v1.3, Qwen-1.5-0.5B-Chat Bai et al. (2023), and Vicuna-13B-v1.5, referred to as Llama-2, Vicuna-7B, Qwen, and Vicuna-13B. Llama-2 is specifically trained for safety alignment, while Vicuna-7B is fine-tuned from Llama without additional safety alignment. Both Llama-2 and Vicuna-7B have 32 layers. Qwen and Vicuna-13B have 24 and 40 layers, respectively. The LLM configurations are detailed in Appendix F.1.
@@
For jailbreak methods, we considered four widely applied ones: optimization-based methods include GCG-Individual and AutoDAN, which iteratively optimize a jailbreak prompt aiming at generating affirmative responses. Model-based jailbreak, i.e., PAIR, uses an attacker LLM to refine the jailbreak prompts. The rule-based method includes DeepInception Li et al. (2023), which crafts jailbreak prompts based on a stealthy template. A detailed illustration of jailbreaks can be found in F.2. Correspondingly, we utilized different defense methods as baselines for ABD: Paraphrase Jain et al. (2023), Retokenization Jain et al. (2023) and Self-Reminder Xie et al. (2023) reformulate the input to avoid attack; PPL Alon and Kamfonas (2023), Self-ExamPhute et al. (2023) and Intention Analysis (IA) Zhang et al. (2025) defense LLMs by double-checking their outputs; SafeDecoding Xu et al. (2024) uses a tuned model to modify the output probability distribution. Further introduction about these defenses can be found in Appendix F.3.
@@
Following Xu et al. (2024), we adopted Just-Eval Lin et al. (2023) to measure the general abilities of LLMs applied defense. Just-Eval is a comprehensive benchmark containing 1,000 diverse instructions, covering seven task types (e.g., reasoning, math, coding, etc.) and seven topics (e.g., ethics, nature, STEM, etc.). Furthermore, following Xu et al. (2024) and Liu et al. (2024d), we utilized the 50-behavior subset from AdvBench Zou et al. (2023) as the test set. This subset, curated by Chao et al. (2023), was created by removing duplicate harmful behavior prompts from AdvBench, ensuring efficiency while mitigating biases caused by repeated behaviors. We used DSR when evaluating defense effectiveness. Following Lin et al. (2023), we leveraged GPT-4o-mini OpenAI (2024) to score the quality of the outputs, ranging from 1 to 5, across five aspects: helpfulness, clarity, factuality, depth, and engagement. We reported the average score for each aspect and their macro-average score, denoted as Avg. To evaluate the efficiency, we reported Runtime per Query, which represents the average time taken to process a single query. We also utilized the above metrics to derive an overall score that simultaneously reflects LLMs’ responding speed and quality. The calculation of the overall score is shown in Appendix F.4.2.
@@
We randomly filtered 400 non-overlapping AdvBench samples to compute $\mu_{\mathcal{D}}^{l}$ and used them as a validation set. We used GCG-Universal Zou et al. (2023) to attack the validation set. GCG-Universal finds a shared jailbreak suffix for all harmful prompts that can deceive LLMs. We set $w=0.8$ when calculating $\mathcal{L}_{\mathrm{Robust}}(\Theta,M)$ as it best balances performances on both attacked and unattached data. We conduct ablation studies on $w$ in Appendix C.3, results shown in Table 5. Appendix E.3 presents settings of ABD optimization.
@@ 5.2 Experimental Results
In all models, the selected layers under the mask $M$ mainly include low and middle layers, such as layers 5 and 12 in Vicuna-7B, layers 2 and 12 in Llama-2, layers 2, 11, and 14 in Vicuna-13B, and layers 5 and 14 in Qwen. These layers are most frequently penalized, consistent with our observations and proposed explanation that low and middle layers are more vulnerable to jailbreak attacks due to significant activation shifts.
@@
Figure 3(c) shows the inclusion ratio, representing the percentage of jailbreak activations within the harmful activation space. The detailed calculation of inclusion ratios is provided in Appendix C.2. Before ABD, the inclusion ratio of various jailbreak activations is below 0.4, indicating that these activations lie outside the safety boundary. After ABD, the inclusion ratio rises to 1, demonstrating ABD’s effectiveness in constraining jailbreak activations. For visualization, Figure 1 projects activations for different prompts. Under ABD, jailbreak activations are progressively constrained within the harmful activation space. This further confirms that ABD progressively confines jailbreak activations within the harmful activation space and mitigates jailbreak effects.
Robust Defense Performance of ABD.
We report the statistical defense results of ABD on the test set in Table 1. We also report representative results on Qwen and Vicuna-13B in Table 2, with full results shown in Table 9 in Appendix G.2.
For Vicuna-7B which lacks specific safety alignment, the defense poses a more challenging task. Nevertheless, ABD demonstrates competitive performance. For jailbreak methods such as PAIR and DeepInception, ABD achieves a higher DSR (58%) compared to Paraphrase (2%) and Retokenization (0%), both of which require costly prompt reformulation. Against the GCG-Individual attack, ABD successfully defends against all jailbreak samples. For the well-aligned model Llama-2-7B-chat, ABD achieves 100% DSR under most jailbreak methods.
For Qwen and Vicuna-13B, ABD maintains nearly 100% DSR, showcasing competitive performance. Notably, for smaller models like Qwen, IA struggles to match its effectiveness on 7B or 13B models, while ABD remains consistently effective. To further validate ABD, we tested it on the full AdvBench and a broader set of jailbreaks. Results are shown in Table 8 in Appendix G.
Efficiency and minimal overhead of ABD.
Table 3 shows Runtime per Query, Just-Eval scores, and overall scores of Vicuna-7B with different defenses applied. We find that ABD has the greatest overall score across all defenses. Moreover, ABD only adds marginal extra time cost and general ability affection. Specifically, it only causes less than $1\%$ delay in each sample and less than $2\%$ perturbation in overall ability, compared to costly methods such as IA, Paraphrase and Self-Exam. Furthermore, with ABD applied, the helpfulness, actuality, depth, and engagement also show a slight increase. We further discover that for baseline defenses such as Retokenization and Self-Reminder, LLM would generate overly simplistic outputs, which leads to smaller Runtime per Query, but they have significantly smaller Just-Eval scores.We also compared defenses on complexity, overhead, and implementation difficulty, highlighting ABD’s efficiency (see Table 10 in Appendix E.5).
@@ 6 Conclusion and Future Work
We conducted a comprehensive study of jailbreak mechanisms, analyzing over 30,000 samples. Our findings reveal that jailbreak shifts harmful activations outside the safety boundary in each layer, with the most severe shifts in the low and middle layers. Motivated by this finding, we proposed ABD, which drives jailbreak activations back within the safety boundary, utilizing LLMs’ intrinsic sensitivity to harmful information. Our experiments suggest that ABD is both practical and efficient. In the near future, we aim to investigate the challenges of jailbreaks in multi-turn dialogue scenarios.
@@
For certain attack methods, under-aligned models (Vicuna-7B-v1.3) may not perform significantly as well-aligned models (Llama2-7B-chat). We believe this is because ABD’s effectiveness depends on safety alignment. Under-aligned models have unclear safety boundaries, which complicate the search for penalty functions that balance general ability and DSR. Therefore, they typically need more extensive optimization. Future work could refine ABD by specifically searching for activation spaces that preserve the concept of "safety", therefore enhancing its generalizability on uncensored models.
@@
In this study, we primarily focus on single-round jailbreak scenarios. We do not extend our analysis to more complex jailbreaks that involve long contexts or multi-round dialogues, such as CFA Sun et al. (2024). As a result, the relationship between jailbreaks and safety boundaries remains largely unexplored. While multi-turn jailbreaks are beyond this paper’s scope, ABD can extend to them by using prior context tokens to compute $\mu^{l}_{\mathcal{D}}$ and applying the same optimization strategy. To adapt ABD for multi-turn jailbreaks, potential improvements include: (1) Dynamic adjustment: Relax or strengthen ABD constraints based on the safety of prior turns. (2) Selective application: Apply ABD at critical points, e.g., when malicious topics emerge, to balance efficiency and effectiveness. (3) Constraint propagation: Carry over constraints across turns to prevent harmful activations from accumulating.
@@
The aim of this research is to enhance the explainability and safety of LLMs. Our proposed jailbreak mechanism, that jailbreak shift activations out of the safety boundary, can mitigate disputes on how jailbreak happens and promote the development of both LLM explainability and safety. We highlight that the development of ABD only needs publicly available datasets and jailbreak methods and does not require designing new jailbreak methods. We demonstrate some harmful responses from LLMs only for illustration. We acknowledge that ABD would cause the development of new attacks. We will explore using random perturbation in the input sequence rather than a particular jailbreak method when optimizing to mitigate such attacks.